I need help with finding the value of 12 to the power 3. (Exponents)
Answer by jim_thompson5910(35256)

12^3 = 12*12*12 = 1728

When you cube something, you are multiplying that number by itself (3 copies total). Other examples

5^3 = 5*5*5 = 125

8^3 = 8*8*8 = 512